A sample of nitrogen gas expands in volume from 1.6 L to 5.4 L at constant temperature. Calculate the work done in joules if the gas expands against a constant pressure of 3.7 atm.


1
Expert's answer
2021-09-18T23:58:16-0400

(c) Work done , W = PΔV

P = 3.7 atm= 3.7× 101.3 ×10^3Pa ( Since 1atm = 101.3 ×10^3 Pa)

= 374810 Pa

So W = 374810 Pa ×3.8 ×10^-3 m^3

=1424.28 J